OpenStudy (anonymous):

x^5-3x^4-15x^3+45x^2-16x+48=0

hero (hero):

\[x^5-3x^4-15x^3+45x^2-16x+48=0 \\x^4(x - 3)-15x^2(x - 3)-16(x - 3) = 0 \\(x - 3)(x^4 - 15x^2 - 16) = 0 \\(x - 3)(x^4 - (16 - 1)x^2 - 16) = 0 \\(x - 3)(x^4 - 16x^2 + x^2 - 16) = 0 \\(x - 3)(x^2(x^2 - 16) + 1(x^2 - 16)) = 0 \\(x - 3)(x^2 - 16)(x^2 + 1) = 0 \\(x - 3)(x + 4)(x - 4)(x^2 + 1) = 0\]